Two vertices of a triangle are two vertex of trinagle are (1,-5) and (5,1) The medians intersect at (1,-1).Find the third vertex of the triangle.
Answer by dkppathak(439)

two vertex of triangle are (1,-5) and (5,1) median intersect at (1, -1)
let third vertex of triangle is (x,y)
mid point side (1,-5) and (5,1) where median meet is 5+1/2=3 and -5+1/2=-2
medians of triangle intersect at the ratio of 2:1
by using section formula
we can say 1= 2 multiplied 3+x/3
3=6+x
x=-3
similarly
-1=-2 multiplied -2+y/3
-3=-4+y
y=1
third vertex will be (-3,1)
answer (-3,1)